# Rethinking Cross-Domain Sequential Recommendation under Open-World Assumptions

### 20 Dec 2023 v2

@@ Abstract.
Cross-Domain Sequential Recommendation (CDSR) methods aim to tackle the data sparsity and cold-start problems present in Single-Domain Sequential Recommendation (SDSR). Existing CDSR works design their elaborate structures relying on overlapping users to propagate the cross-domain information. However, current CDSR methods make closed-world assumptions, assuming fully overlapping users across multiple domains and that the data distribution remains unchanged from the training environment to the test environment. As a result, these methods typically result in lower performance on online real-world platforms due to the data distribution shifts. To address these challenges under open-world assumptions, we design an Adaptive Multi-Interest Debiasing framework for cross-domain sequential recommendation (AMID), which consists of a multi-interest information module (MIM) and a doubly robust estimator (DRE). Our framework is adaptive for open-world environments and can improve the model of most off-the-shelf single-domain sequential backbone models for CDSR. Our MIM establishes interest groups that consider both overlapping and non-overlapping users, allowing us to effectively explore user intent and explicit interest. To alleviate biases across multiple domains, we developed the DRE for the CDSR methods. We also provide a theoretical analysis that demonstrates the superiority of our proposed estimator in terms of bias and tail bound, compared to the IPS estimator used in previous work. To promote related research in the community under open-world assumptions, we collected an industry financial CDSR dataset from Alipay, called "MYbank-CDR". Extensive offline experiments on four industry CDSR scenarios including the Amazon and MYbank-CDR datasets demonstrate the remarkable performance of our proposed approach. Additionally, we conducted a standard A/B test on Alipay, a large-scale financial platform with over one billion users, to validate the effectiveness of our model under open-world assumptions.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; ; price: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 242ccs: Information systems  Recommender systemsccs: Computing methodologies Neural networks
@@ 1. Introduction
Single-domain sequential recommendation (SDSR) has garnered wide attention in E-commerce recommender systems, for its ability to model dynamic user preferences based on their sequential interactions. Nevertheless, traditional sequential recommendation methods are often limited by data sparsity and the cold-start problem, affecting their recommendation performance (Sun et al., 2019; Hidasi et al., 2015; Kang and McAuley, 2018).
@@
To alleviate these problems, several Cross Domain Sequential Recommendation (CDSR) approaches (Hu et al., 2018; Ma et al., 2019; Ouyang et al., 2020) have been proposed to transfer the knowledge between the source domain and target domain. Existing CDSR approaches commonly employ feature combination (Zhu et al., 2020; Zhang et al., 2018; Ma et al., 2019) or bi-directional transfer mapping strategies (Li and Tuzhilin, 2020; Li et al., 2021a; Zhao et al., 2017) to enhance recommendation accuracy by leveraging the overlapping users. Most of these methods conduct experiments under closed-world assumptions, where multiple domains share fully overlapping users and data distribution remains unchanged from the training environment to the test environment. However, such closed-world assumptions do not exist in industry recommendation platforms, which operate under open-world assumptions instead. In this study, the concept of the “open-world is used in contrast to the closed world. The term "open-world" (Wu et al., 2021; Yue et al., 2020) is employed to describe and summarize the scenario, which contains minority overlapping users among multiple domains and contains the selection bias in modeling. In the open-world environment, the input/output space expands, and data distribution shifts due to unseen factors. Therefore, to enhance the model’s performance in CDSR, it is imperative to address the challenges that arise under open-world assumptions.
@@
First of all, for multiple domains with few overlapping users, how to construct the model with the non-overlapping users and improve the recommendation performance? Most previous CDSR methods cannot be directly extended to handle partially overlapping CDSR settings, especially when there are only a few shared users across domains that are often encountered in the real world. To address this challenge, recent researches (Li and Tuzhilin, 2020; Cui et al., 2020; Guo et al., 2021; Ma et al., 2022) have explored the use of graphic deep learning to enhance both overlapping and non-overlapping user embeddings and propagate interest information among non-overlapping users. More recently, Cao et al. (Cao et al., 2022a) have designed a graph neural network with a contrastive cross-domain infomax objective to improve the learning of non-overlapping user embeddings. Nevertheless, in partially overlapping CDSR scenarios, such methods have great limitations and will drop their performance under open-world assumptions. These CDSR approaches still rely heavily on overlapping users, with more than 70% of users being common across domains, to establish bridge connections among multiple domains and perform knowledge aggregation and transition processes.
@@
2. Motivation: Towards Open-World CDSR
@@ 3.1. Problem Definition
In this paper, we consider a partially overlapping CDSR scenario composed of multiple domains $\mathcal{Z}=\{Z_{1},...,Z_{|\mathcal{Z}|}\}$. Let $\mathcal{U}=\{u_{1},...,$ $u_{|\mathcal{U}|}\}$, $\mathcal{V}=\{v_{1}^{Z_{1}},...,v_{|\mathcal{V}|}^{Z_{|\mathcal{Z}|}}\}$ be the user set, item set, rating set. A user who only has historical behaviors in one domain is referred to as a non-overlapping user, while a user with historical behaviors in multiple domains is referred to as an overlapping user. As a certain user, denote $\mathcal{S}=\{S^{Z_{1}},...,S^{Z_{|\mathcal{Z}|}})\}$ the corresponding sequential behaviours of the users. For example, $S^{Z_{1}}=\{v_{1},...,v_{T}\}$ represents the single-domain sequence, where $T$ is a variable length. Given the data $\mathcal{D}=\mathcal{U}\times\mathcal{V}$, CDSR aims to develop a personalized ranking function that utilizes the past item sequences from multiple domains of a user and predicts the next item (i.e. $v_{T+1}$) in each domain that the user is most likely to choose.
@@
where $r_{u,v}$ denotes the prediction from user $u$ to item $v$ in domain $X$. However, the objective of CDSR approaches is to predict the next item for a given user $u$ based on their interaction sequences:
(2) $\mathrm{argmax}\;\;P^{X}\left(v^{X}_{|S^{X}|+1}=v|S^{X},S^{Y},U^{X},U^{Y},V^{X}\right),\mathrm{if}\;\;v\in\mathcal{V}^{X}.\$
@@ 4.3. Doubly Robust Estimator for CDSR
where $e^{Z}_{u,v}=|\hat{r}^{Z}_{u,v}-r^{Z}_{u,v}|$ or $e^{Z}_{u,v}=(\hat{r}^{Z}_{u,v}-r^{Z}_{u,v})^{2}$ via optional measure metrics for MAE or MSE. The imputation error $\hat{e}^{Z}_{u,v}=g_{\phi^{Z}}(\mathrm{Mean}(\bm{S}_{u}^{*Z})||\mathbf{v}^{Z}))$ is computed by imputation model which aims to estimate the prediction error $e_{u,v}$ on the observed data. We also learn the propensity $\hat{p}_{u,v}=g_{\psi^{Z}}(\mathrm{Mean}(\bm{S}_{u}^{*Z})||\mathbf{v}^{Z}))$. The imputation model $g_{\phi^{Z}}$ and the propensity model $g_{\psi^{Z}}$ are implemented in a multi-task manner. The bias of the estimator is derived as follows.
@@ 5.1. Experimental Setup
To conduct our experiments, we followed the methodology of previous research (Xu et al., 2023; Cao et al., 2022a) and used the Amazon datasets  [^3] http://jmcauley.ucsd.edu/data/amazon/index 2014.html, which consist of 24 item domains. We selected two pairs of domains, namely “Cloth-Sport" and “Phone-Elec", and formulated two different tasks. Besides, we collect a Table 3 summarizes the statistics for each task. To replicate selection bias in online platforms, we adjusted the non-overlapping ratio $\mathcal{K}_{u}$ for each dataset, selecting from $\{25\%,75\%\}$. Varying the ratio results in different numbers of non-overlapping users across domains, with a higher $\mathcal{K}_{u}$ indicating a less biased environment. Following previous CDSR literature (Ma et al., 2019; Li et al., 2021a; Cao et al., 2022a), we employed the leave-one-out technique to evaluate the performance of our developed model. In order to ensure an unbiased evaluation, we adopt the methodology employed in previous works (Krichene and Rendle, 2020; Zhao et al., 2020), wherein we randomly select 999 negative items (i.e., items that the user has not interacted with) and combine them with 1 positive item (i.e., a ground-truth interaction) to form our recommendation candidates for the ranking test. We evaluated our model using the normalized discounted cumulative gain (NDCG@10) and hit rate (HR@10) metrics, which are common in CDSR literature. For all comparative models, we ran each experiment five times and reported the results by mean and variance. Further details about our experimental setup are in Appendix C.1.
@@ 5.3. Online A/B Test
5.4. Hyperparameter Analysis
The threshold $k$ for the group. We conduct ablation experiments by varying the threshold $k\in\{0.5,0.6,0.7,0.8,0.9\}$ to investigate the impact of the threshold $k$ on constructing interest groups in the multi-interest information module. Our results show that a larger threshold $k$ (0.5 $\rightarrow$ 0.7) leads to better performance, as more related interest information can be transferred. However, when the threshold $k$ is increased beyond 0.7, the model’s performance drops due to noise interference and redundant information. Therefore, to achieve superior performance, we set the threshold $k$ to 0.7.
The number of the sampled users. To explore the impact of the number of sampled users on the multi-interest information module, we conduct ablation experiments varying the number of sampled users from 128 to 1024. Our findings suggest that an increase in the number of sampled users initially improves the recommendation performance, but it eventually declines when the matching neighbors reach 1024. This observation indicates that having too few sampled users would provide insufficient transferred information, while too many sampled users could introduce interference noise and compromise the model’s performance. In practice, we select the number of sampled users to be 512 as it results in the best performance for our model. More analysis and results can be found in Appendix C.3.
(a) Cloth Domain
(b) Sport Domain
(c) Cloth Domain
(d) Sport Domain
Figure 6. Impact of the threshold $k$ and the number of the sampled users.
@@ 6. Related Work
Conventional cross-domain recommendation has emerged as a promising solution for mitigating data sparsity and cold-start issues encountered in single-domain recommendation systems. Early CDR studies (Hu et al., 2018; Liu et al., 2020b) have primarily focused on developing approaches that transfer cross-domain knowledge by relying on overlapping users. However, real-world CDR scenarios often do not satisfy strict overlapping requirements and exhibit only a small fraction of common users across domains. To tackle this challenge, recent methods (Sheng, 2021; Luo et al., 2023)) have proposed a network structure comprising shared and domain-specific networks to effectively capture the unique characteristics and commonalities across all domains simultaneously. While these CDR approaches incorporate valuable information from relevant domains to enhance performance in the target domain, they still encounter difficulties in addressing the contextual sequential dependencies within users’ interaction history, which are essential for comprehensive modeling in CDSR tasks.
@@ 7. Conclusions and Discussions
Limitations. Our MIM constructs interest groups between pairs of domains, which share cross-domain knowledge as widely as possible. In the real-world platform, commercial activities are usually composed of multiple domains. However, constructing all the groups for $|\mathcal{Z}|$ domains has a time complexity of $\mathcal{O}(|\mathcal{Z}|^{2})$, which can become quite time-consuming as the number of domains increases. Therefore, it is important to develop more efficient methods for constructing groups among multiple domains in future work.
@@ Acknowledgements.
This work is supported by MYbank, Ant Group.
@@ C.1. Experiment Setup
Evaluation Metrics In each domain, we divided the users into three sets: 80% for training, 10% for validation, and 10% for testing. To preprocess the data, items with fewer than 10 interactions and users with fewer than 5 interactions in their respective domains were filtered out, following the approach of previous studies (Xu et al., 2023; Cao et al., 2022b). This ensured that the embeddings learned by the users/items were representative of their source domain. A non-overlapping ratio $\mathcal{K}_{u}$ was introduced to control the number of non-overlapping users and simulate different debiased scenarios. For example, in the Amazon "Cloth-Sport" dataset with $\mathcal{K}_{u}$ = 25%, the number of overlapped users in the training set was calculated as (27,519 + 107,984 - 16,377 * 2) * 0.25 * 0.8 = 20,549. The same sampling strategy was applied to the validation set, while the test set was not downsampled. This sampling strategy can simulate the occurrence of selection bias in the open-world environment, where the training set mainly consists of users who are more likely to be selected or exposed, while the test set covers a wider range of users without careful selection. All evaluation metrics used in this study indicate better performance with higher values. Regarding the unseen users in $\mathcal{D}$, we use the non-overlapping users who were not selected for the observed data as a substitute for unseen users. We remove the actual ratings for the unseen users while retaining their sequences.
@@ C.3. Hyperparameter Analysis
Appendix D Potential Societal Impacts

### 29 Feb 2024 v5

@@ Abstract.
Cross-Domain Sequential Recommendation (CDSR) methods aim to tackle the data sparsity and cold-start problems present in Single-Domain Sequential Recommendation (SDSR). Existing CDSR works design their elaborate structures relying on overlapping users to propagate the cross-domain information. However, current CDSR methods make closed-world assumptions, assuming fully overlapping users across multiple domains and that the data distribution remains unchanged from the training environment to the test environment. As a result, these methods typically result in lower performance on online real-world platforms due to the data distribution shifts. To address these challenges under open-world assumptions, we design an Adaptive Multi-Interest Debiasing framework for cross-domain sequential recommendation (AMID), which consists of a multi-interest information module (MIM) and a doubly robust estimator (DRE). Our framework is adaptive for open-world environments and can improve the model of most off-the-shelf single-domain sequential backbone models for CDSR. Our MIM establishes interest groups that consider both overlapping and non-overlapping users, allowing us to effectively explore user intent and explicit interest. To alleviate biases across multiple domains, we developed the DRE for the CDSR methods. We also provide a theoretical analysis that demonstrates the superiority of our proposed estimator in terms of bias and tail bound, compared to the IPS estimator used in previous work. To promote related research in the community under open-world assumptions, we collected an industry financial CDSR dataset from Alipay, called "MYbank-CDR". Extensive offline experiments on four industry CDSR scenarios including the Amazon and MYbank-CDR datasets demonstrate the remarkable performance of our proposed approach. Additionally, we conducted a standard A/B test on Alipay, a large-scale financial platform with over one billion users, to validate the effectiveness of our model under open-world assumptions. Code and dataset are available at https://github.com/WujiangXu/AMID.
@@
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3589334.3645351copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporeisbn: 979-8-4007-0171-9/24/05submissionid: 242ccs: Information systems  Recommender systemsccs: Computing methodologies Neural networks
@@ 1. Introduction
Single-domain sequential recommendation (SDSR) has garnered wide attention in E-commerce recommender systems, for its ability to model dynamic user preferences based on their sequential interactions. Nevertheless, traditional sequential recommendation methods are often limited by data sparsity and the cold-start problem, affecting their recommendation performance (Sun et al., 2019; Hidasi et al., 2015; Kang and McAuley, 2018; Wu et al., 2019; Wang et al., 2022).
@@
To alleviate these problems, several Cross Domain Sequential Recommendation (CDSR) approaches (Hu et al., 2018; Ma et al., 2019; Ouyang et al., 2020) have been proposed to transfer the knowledge between the source domain and target domain. Existing CDSR approaches commonly employ feature combination (Zhu et al., 2020; Zhang et al., 2018; Ma et al., 2019) or bi-directional transfer mapping strategies (Li and Tuzhilin, 2020; Li et al., 2021a; Zhao et al., 2017) to enhance recommendation accuracy by leveraging the overlapping users. Most of these methods conduct experiments under closed-world assumptions, where multiple domains share fully overlapping users and data distribution remains unchanged from the training environment to the test environment. However, such closed-world assumptions do not exist in industry recommendation platforms, which operate under open-world assumptions instead. In this study, the concept of the “open-world is used in contrast to the closed world. The term "open-world" (Wu et al., 2021b, a; Yue et al., 2020) is employed to describe and summarize the scenario, which contains minority overlapping users among multiple domains and contains the selection bias in modeling. In the open-world environment, the input/output space expands, and data distribution shifts due to unseen factors. Therefore, to enhance the model’s performance in CDSR, it is imperative to address the challenges that arise under open-world assumptions.
@@
First of all, for multiple domains with few overlapping users, how to construct the model with the non-overlapping users and improve the recommendation performance? Most previous CDSR methods cannot be directly extended to handle partially overlapping CDSR settings, especially when there are only a few shared users across domains that are often encountered in the real world. To address this challenge, recent researches (Li and Tuzhilin, 2020; Cui et al., 2020; Guo et al., 2021b; Ma et al., 2022) have explored the use of graph deep learning to enhance both overlapping and non-overlapping user embeddings and propagate interest information among non-overlapping users. More recently, Cao et al. (Cao et al., 2022a) have designed a graph neural network with a contrastive cross-domain infomax objective to improve the learning of non-overlapping user embeddings. Nevertheless, in partially overlapping CDSR scenarios, such methods have great limitations and will drop their performance under open-world assumptions. These CDSR approaches still rely heavily on overlapping users, with more than 70% of users being common across domains, to establish bridge connections among multiple domains and perform knowledge aggregation and transition processes.
@@
2. Motivation: Towards Open-World Cross-Domain Recommendation
@@ 3.1. Problem Definition
In this paper, we consider a partially overlapping CDSR scenario composed of multiple domains $\mathcal{Z}=\{Z_{1},...,Z_{|\mathcal{Z}|}\}$. Let $\mathcal{U}=\{u_{1},...,$ $u_{|\mathcal{U}|}\}$, $\mathcal{V}=\{v_{1}^{Z_{1}},...,v_{|\mathcal{V}|}^{Z_{|\mathcal{Z}|}}\}$ be the user set, item set. A user who only has historical behaviors in one domain is referred to as a non-overlapping user, while a user with historical behaviors in multiple domains is referred to as an overlapping user. As a certain user, denote $\mathcal{S}=\{S^{Z_{1}},...,S^{Z_{|\mathcal{Z}|}})\}$ the corresponding sequential behaviours of the users. For example, $S^{Z_{1}}=\{v_{1},...,v_{T}\}$ represents the single-domain sequence, where $T$ is a variable length. Given the data $\mathcal{D}=\mathcal{U}\times\mathcal{V}$, CDSR aims to develop a personalized ranking function that utilizes the past item sequences from multiple domains of a user and predicts the next item (i.e. $v_{T+1}$) in each domain that the user is most likely to choose.
@@
where $r_{u,v}$ denotes the prediction from user $u$ to item $v$ in domain $X$. $Y$ represents a different domain. However, the objective of CDSR approaches is to predict the next item for a given user $u$ based on their interaction sequences:
(2) $\mathrm{argmax}\;\;P^{X}\left(S^{X}_{|S^{X}|+1}=v|S^{X},S^{Y},U^{X},U^{Y},V^{X}\right),\mathrm{if}\;\;v\in\mathcal{V}^{X}.\$
@@ 4.3. Doubly Robust Estimator for CDSR
where $e^{Z}_{u,v}=|\hat{r}^{Z}_{u,v}-r^{Z}_{u,v}|$ or $e^{Z}_{u,v}=(\hat{r}^{Z}_{u,v}-r^{Z}_{u,v})^{2}$ via optional measure metrics for MAE or MSE. $o$ denotes the observation indicator, signifying whether a user-item sample is observed. The imputation error $\hat{e}^{Z}_{u,v}=g_{\phi^{Z}}(\mathrm{Mean}(\bm{S}_{u}^{*Z})||\mathbf{v}^{Z}))$ is computed by imputation model which aims to estimate the prediction error $e_{u,v}$ on the observed data. We also learn the propensity $\hat{p}_{u,v}=g_{\psi^{Z}}(\mathrm{Mean}(\bm{S}_{u}^{*Z})||\mathbf{v}^{Z}))$. The imputation model $g_{\phi^{Z}}$ and the propensity model $g_{\psi^{Z}}$ are implemented in a multi-task manner. The bias of the estimator is derived as follows.
@@ 5.1. Experimental Setup
To conduct our experiments, we followed the methodology of previous research (Xu et al., 2023a; Cao et al., 2022a) and used the Amazon datasets  [^3] http://jmcauley.ucsd.edu/data/amazon/index 2014.html, which consist of 24 item domains. We selected two pairs of domains, namely “Cloth-Sport" and “Phone-Elec", and formulated two different tasks. Besides, we collect a Table 3 summarizes the statistics for each task. To replicate selection bias in online platforms, we adjusted the non-overlapping ratio $\mathcal{K}_{u}$ for each dataset, selecting from $\{25\%,75\%\}$. Varying the ratio results in different numbers of non-overlapping users across domains, with a higher $\mathcal{K}_{u}$ indicating a less biased environment. In order to ensure an unbiased evaluation, we adopt the methodology employed in previous works (Krichene and Rendle, 2020; Zhao et al., 2020), wherein we randomly select 999 negative items (i.e., items that the user has not interacted with) and combine them with 1 positive item (i.e., a ground-truth interaction) to form our recommendation candidates for the ranking test. We evaluated our model using the normalized discounted cumulative gain (NDCG@10) and hit rate (HR@10) metrics, which are common in CDSR literature. For all comparative models, we ran each experiment five times and reported the results by mean and variance. Further details about our experimental setup are in Appendix C.1.
@@ 6. Related Work
Conventional cross-domain recommendation has emerged as a promising solution for mitigating data sparsity and cold-start issues encountered in single-domain recommendation systems. Early CDR studies (Hu et al., 2018; Liu et al., 2020b; Yuan et al., 2020, 2021) have primarily focused on developing approaches that transfer cross-domain knowledge by relying on overlapping users. However, real-world CDR scenarios often do not satisfy strict overlapping requirements and exhibit only a small fraction of common users across domains. To tackle this challenge, recent methods (Sheng, 2021; Luo et al., 2023; Xu et al., 2023b; Guo et al., 2023a; Wu et al., 2022, 2023) have proposed a network structure comprising shared and domain-specific networks to effectively capture the unique characteristics and commonalities across all domains simultaneously. While these CDR approaches incorporate valuable information from relevant domains to enhance performance in the target domain, they still encounter difficulties in addressing the contextual sequential dependencies within users’ interaction history, which are essential for comprehensive modeling in CDSR tasks.
@@
Multi-interest recommendation Recent sequential recommendation models (Chen et al., 2020, 2021; Cen et al., 2020) increasingly focus on multi-interest representation learning to capture diverse user interests, akin to disentangled representation. For instance, DIN (Zhou et al., 2018) uses historical user behavior to discern interests, while DIEN (Zhou et al., 2019) enhances this with a GRU and attentional update gate to model evolving interests over time. However, their performance may be limited by extracting a single typical interest. To address this, current models employ multi-interest extraction modules. MIND (Li et al., 2019) uses capsule routing and label-aware attention to capture diverse interests during matching. M5 (Zhao et al., 2023) learns coarse-grained and fine-grained interests from both ID and content graph embeddings. IESRec (Liu et al., 2023) exploits user preferences via a temporal optimal transport algorithm in non-overlapped cross-domain recommendation. Unlike existing methods, our approach is simple yet effective and transfers cross-domain information for all users without relying on side information like text reviews or product features.
@@ Acknowledgements.
This work is partly supported by MYbank, Ant Group, NSFC (62222607) and STCSM (22511105100).
@@ C.1. Experiment Setup
Evaluation Metrics In each domain, we divided the users into three sets: 80% for training, 10% for validation, and 10% for testing. To preprocess the data, items with fewer than 10 interactions and users with fewer than 5 interactions in their respective domains were filtered out, following the approach of previous studies (Xu et al., 2023a; Cao et al., 2022b). This ensured that the embeddings learned by the users/items were representative of their source domain. A non-overlapping ratio $\mathcal{K}_{u}$ was introduced to control the number of non-overlapping users and simulate different debiased scenarios. For example, in the Amazon "Cloth-Sport" dataset with $\mathcal{K}_{u}$ = 25%, the number of overlapped users in the training set was calculated as (27,519 + 107,984 - 16,377 * 2) * 0.25 * 0.8 = 20,549. The same sampling strategy was applied to the validation set, while the test set was not downsampled. This sampling strategy can simulate the occurrence of selection bias in the open-world environment, where the training set mainly consists of users who are more likely to be selected or exposed, while the test set covers a wider range of users without careful selection. All evaluation metrics used in this study indicate better performance with higher values. Regarding the unseen users in $\mathcal{D}$, we use the non-overlapping users who were not selected for the observed data as a substitute for unseen users. We remove the actual ratings for the unseen users while retaining their sequences.
@@ C.3. Hyperparameter Analysis
Appendix D Limitations
Our MIM constructs interest groups between pairs of domains, which share cross-domain knowledge as widely as possible. In the real-world platform, commercial activities are usually composed of multiple domains. However, constructing all the groups for $|\mathcal{Z}|$ domains has a time complexity of $\mathcal{O}(|\mathcal{Z}|^{2})$, which can become quite time-consuming as the number of domains increases. Therefore, it is important to develop more efficient methods for constructing groups among multiple domains in future work.
Appendix E Potential Societal Impacts